Exercise 7-21

Presented below is information for Metlock Company.
1. Beginning-of-the-year Accounts Receivable balance was $23,200.
2. Net sales (all on account) for the year were $106,000. Metlock does not offer cash discounts.
3. Collections on accounts receivable during the year were $82,500.

Metlock is planning to factor some accounts receivable at the end of the year. Accounts totaling $12,400 will be transferred to Credit Factors, Inc. with recourse. Credit Factors will retain 7% of the balances for probable adjustments and assesses a finance charge of 6%. The fair value of the recourse obligation is $1,188.
Prepare the journal entry to record the sale of the receivables. (If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when the amount is entered. Do not indent manually.)

Account Titles and Explanation

Debit

Credit

SHOW LIST OF ACCOUNTS

Accounts Payable
Accounts Receivable
Advertising Expense
Allowance for Doubtful Accounts
Bad Debt Expense
Cash
Cash Over and Short
Due from Factor
Due to Customer
Discount on Notes Payable
Discount on Notes Receivable
Freight-In
Freight-Out
Gain on Disposal of Land
Interest Expense
Interest Payable
Interest Receivable
Interest Revenue
Inventory
Land
Loss on Sale of Receivables
Miscellaneous Expenses
No Entry
Notes Payable
Notes Receivable
Office Expense
Owner's Drawings
Petty Cash
Postage Expense
Purchase Discounts
Recourse Liability
Sales Discounts
Sales Discounts Forfeited
Sales Revenue
Service Revenue
Supplies
Supplies Expenses
Unearned Sales Revenue

Compute Metlock’s accounts receivable turnover for the year, assuming the receivables are sold. (Round answers to 2 decimal places, e.g. 4.57.)
Accounts receivable turnover

times
Days to collect accounts receivable

days
Answer #1
Journal Entry-metlock Company
S. No. Accounts Titles and Explanations Debit Credit
1 No entry
2 Accounts Receivables 106,000
Sales 106,000
To Record Sales made on Account
3 Cash 82,500
Accounts Receivables 82,500
To Record Collection Made
Acccount receivable Turnover=Credit sales / Average receivable
Average receivable = (Beginning Receivable + Ending Receivable)/2
( 23200+ ( 23200+106000-82500))/2
$34950/2=17475
Account Receivable Turnover= $106000/17475=6.07 times
Day to Collect account receivable = 365/ Account receivable turnover ratio
=365/6.07=60.13 days